How many inches is a 17- yard rope ?
Found 2 solutions by nellie12, lenny460:
Answer by nellie12(2)

There are 36 inches in 1 yard, therefore there would be 612 inches in a 17 yard rope (36*17=612)

Answer by lenny460(1073)

How many inches is a 17 - yard rope ?
1 yard = 36 inches.
Therefore:
17 yards = 36 * 17 = 612 inches.

Answer:
612 inches.